OpenStudy (anonymous):

A quadrilateral (Not Regular) is inscribed in a circle. One of the angles is 82°, find the angle that is opposite the given angle. Show all calculations.

OpenStudy (anonymous):

the sum of opposite angles of a quadrilateral inscribed in a circle is 180 deg

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

do you want to see a proof ?

OpenStudy (anonymous):

sure :) that would also help me.

OpenStudy (anonymous):

so first we need to use the relation between arc length and angle I hope you know it : arclength = angle * radius ?

OpenStudy (anonymous):

Yes I get that part

OpenStudy (anonymous):

good now :

OpenStudy (anonymous):



OpenStudy (anonymous):

since we measure the angles from the sides of the circle and not from its center the are lengths are multiplied by 2 for each of those angles! so as you can see the angle C can give us arc XYZ length : arclength(XYZ) = 2* C * r and the angle A can give us arc ZWX lenght: arclength(ZWX) = 2* A * r now we can use the fact that the arcs together completes the circle arclength(XYZ) + arclength(ZWX) = 360 * r ( in fact we usually write 2*pi*r) so we have 2*C*r + 2*A*r = 360*r A+C = 180

OpenStudy (anonymous):

Wow thanks I get It You explained it really good. Your amazing :)

OpenStudy (anonymous):

the multiplication by 2 is a bit tricky almost missed it myself lol

OpenStudy (anonymous):

lol :)